### v2

# Fairness Risks for Group-Conditionally Missing Demographics

Abstract
Fairness-aware classification models have gained increasing attention in recent years as concerns grow on discrimination against some demographic groups. Most existing models require full knowledge of the sensitive features, which can be impractical due to privacy, legal issues, and an individual’s fear of discrimination. The key challenge we will address is the group dependency of the unavailability, e.g., people of some race may be more reluctant to reveal their race. Our solution augments general fairness risks with probabilistic imputations of the sensitive features, while jointly learning the group-conditionally missing probabilities in a variational auto-encoder. Our model is demonstrated effective on both image and tabular datasets, achieving an improved balance between accuracy and fairness.
1 Introduction
As machine learning systems rapidly acquire new capabilities and get widely deployed to make human-impacting decisions, ethical concerns such as fairness have recently attracted significant effort in the community. To combat societal bias and discrimination in the model that largely inherit from the training data, a bulk of research efforts have been devoted to addressing group unfairness, where the model performs more favorably (e.g., accurately) to one demographic group than another (Hardt et al., 2016; Dwork et al., 2012; Ye and Xie, 2020; Buolamwini and Gebru, 2018; Gianfrancesco et al., 2018; Mehrabi et al., 2021; Yapo and Weiss, 2018). Group fairness typically concerns both the conventional classification labels (e.g., recidivism) and socially sensitive group features (e.g., gender, or race). As applications generally differ in their context of fairness, a number of quantitative metrics have been developed such as demographic parity, equal opportunity, and equalized odds. A necessarily outdated overview is available at Barocas et al. (2019) and Wikipedia contributors (2024).
Learning algorithms for group fairness can be broadly categorized into pre-, post-, and in-processing methods. Pre-processing methods transform the input data to remove dependence between the class and demographics according to a predefined fairness constraint (Kamiran and Calders, 2012). Post-processing methods warp the class labels (or their distributions) from any classifier to fulfill the desired fairness criteria (Hardt et al., 2016). In-processing approaches, which are most commonly employed including this work, learn to minimize the prediction loss while upholding fairness regularizations simultaneously. More discussions are given in Section 2.
Most of these methods require accessing the sensitive demographic feature, which is often unavailable due to fear of discrimination and social desirability (Krumpal, 2013), privacy concerns, and legal regulations (Coston et al., 2019; Lahoti et al., 2020). For example, people of some gender may be more inclined to withhold their gender information when applying for jobs dominated by other genders. Recently several methods have been developed for this data regime. Lahoti et al. (2020) achieved Rawlsian max-min fairness by leveraging computationally-identifiable errors in adversarial reweighted learning. Hashimoto et al. (2018) proposed a distributionally robust optimization to minimize the risk over the worst-case group distribution. However, they are not effective in group fairness and cannot be customized for different group fairness metrics. Yan et al. (2020) infers groups by clustering the data, but there is no guarantee that the uncovered groups are consistent with the real sensitive features of interest. For example, the former may identify race while we seek to be fair in gender.
We aim to tackle these issues in a complementary setting, where demographics are partially missing. Although their availability is limited, it is still often feasible to obtain a small amount of labeled demographics. [^1] According to Consumer Financial Protection Bureau (2023), “A creditor shall not inquire about the race, color, religion, national origin, or sex of an applicant.” Exceptions include inquiry “for the purpose of conducting a self-test that meets the requirements of §1002.15”, which essentially tests if the fair lending rules are conformed with. For example, some people may not mind disclosing their race or gender. Therefore, different from the aforementioned works that assume complete unavailability of demographics, we study in this paper the semi-supervised setting where they can be partially available at random in both training and test data. Similarly, the class labels can be missing in training.
Semi-supervised learning (SSL) has been well studied, and can be applied to impute missing demographics. Zhang and Long (2021) simply used the labeled subset to estimate the fairness metrics, but their focus is on analyzing its estimation bias instead of using it to train a classifier. Jung et al. (2022) assigned pseudo group labels by training an auxiliary group classifier, and assigned low-confidence samples to random groups. However, the confidence threshold needs to be tuned based on the conditional probability of groups given the class label, which becomes challenging as the latter is also only partially available in our setting.
Dai and Wang (2021) used SSL on graph neural networks to infer the missing demographics, but rounded the predictions to binary group memberships. This is suboptimal because, due to the scarcity of labeled data, there is marked uncertainty in demographic imputations. As a result, the common practice of “rounding” sensitive estimates—so that fairness metrics defined on categorical memberships can be enforced instead of mere independence—may over-commit to a group just by chance, dropping important uncertainty information when the results are fed to downstream learners. Moreover, most applications do not naturally employ an underlying graph, and the method only enforces min-max (i.e., adversarial) fairness instead of group fairness metrics.
Our first contribution, therefore, is to leverage the probabilistic imputation by designing a differentiable fairness risk in Section 4, such that it is customized for a user-specified fairness metric with discrete demographics as opposed to simple independence relationships, and can be directly integrated into general SSL algorithms to regularize the learned posterior towards low risks in both classification and fairness.
Our method is distinguished from generic SSL recipes which predict unobserved protected attributes and then optimize the fairness risk. Performing the two steps separately prevents the common backbone features to be synthesized that concurrently facilitate identifying demographics and class labels. This could be ameliorated via a bi-level optimization, but its computational cost can be high. We hence resort to flattening the two levels into one joint optimization, and then address the pathological phenomenon where the group memberships are also “learned” to promote fairness.
Interestingly, we found an efficient solution by stopping the gradient of fairness risk with respect to the group estimates. Furthermore, to effectively implement this principle, we employed a Monte-Carlo evaluation of the risk, which significantly accelerates inference and differentiation via a provable $O(1/\epsilon^{2})$ sample complexity. So our second contribution hits two birds with one stone, as detailed in Section 4.3.
As our third contribution, we instantiated the semi-supervised classifier with a new encoder and decoder in a variational autoencoder (VAE, Narayanaswamy et al., 2017; Kingma et al., 2014), allowing group-conditional missing demographics, which has so far only been considered for noisy but not missing demographics. In general, the chance of unavailability does depend on specific groups – if people of a certain race are aware of the unfairness against them, they will be more reluctant to disclose their race. The comprehensive model will be introduced in Section 5.
Our fourth contribution is to demonstrate, in Section 6, that our method empirically outperforms state of the art for fair classification where both the demographics and class labels are only partially available. The paper is primed with preliminaries in Section 3.
2 Related Work
Noisy demographics have been tackled by, e.g., Lamy et al. (2019); Celis et al. (2021a). Adversarially perturbed and privatized demographics were studied by Celis et al. (2021b) and Mozannar et al. (2020), which also account for group-conditional noise. However, although they conceptually subsume unavailability as a type of noise, their method and analysis do not carry through. Some of these methods, along with Wang et al. (2020), also require an auxiliary dataset to infer the noise model, which is much harder for missing demographics. Shah et al. (2023) dispenses with such a dataset, but its favorable theoretical properties apply only to Gaussian data, while the bootstrapping-based extension to non-Gaussian data does not model the conditional distribution of demographics.
Another strategy to tackle missing demographics resorts to proxy features (Gupta et al., 2018; Chen et al., 2019; Kallus et al., 2022). Based on domain knowledge, they are assumed to correlate with and allude to the sensitive feature in question. Zhao et al. (2022) optimizes the correlations between these features and the prediction, but it is not tailored to the specific group fairness metric in question. This is difficult because it relies on the binary group memberships which is not modeled by the method. Zhu et al. (2023) considered the accuracy of estimating the fairness metrics using proxies, but did not demonstrate its effectiveness when applied to train a fair classifier. In addition, proxy features can often be difficult to identify. In face recognition where the sensitive feature is age, race, or wearing glasses, raw pixels are no good proxies and identifying semantic features as proxies can be challenging.
3 Preliminary
We first consider supervised learning with group fairness, where samples are drawn from a distribution $D$ over $\mathcal{X}\times\{0,1\}\times\{0,1\}$, with $(X,A,Y)\sim D$. Here $X$ is the non-protected feature, $Y$ is the binary label, and $A$ is the binary sensitive feature which is also referred to as the group demographics. Our method can be extended to multi-class labels and sensitive features, under any fairness risk defined for fully observed $A$ and $Y$. The details are deferred to Appendix A.1. Since our focus is on addressing missing demographics, we will stick with the binary formulation for ease of exposition.
We follow the setup in Menon and Williamson (2018) and aim to find a measurable randomized classifier parameterized by a function $f:\mathcal{X}\to[0,1]$ (e.g., a neural network), such that it predicts an $X\in\mathcal{X}$ to be positive with probability $f(X)$. That is, the predicted label $\hat{Y}\in\{0,1\}$ follows the Bernoulli distribution $\hat{Y}|X\sim\text{Bern}(f(X))$. We denote the conditional distribution as $P_{f}(\hat{Y}|X)$, and omit the subscript $f$ for brevity when the context is clear. This setup also subsumes a classifier based on a class-probability estimator, where $f(X)\in[0,1]$ is translated to a hard deterministic label via a learnable threshold $\tau$: $P_{f}(\hat{Y}=1|X)=[\![f(X)>\tau]\!]$. Here, the Iverson bracket $[\![\cdot]\!]=1$ if $\cdot$ is true, and 0 otherwise.
Suppose we have a training set $\mathcal{D}_{tr}:=\{\mathbf{\bm{x}}_{i},y_{i},a_{i}\}_{i=1}^{n}$ with size $n$, where $\mathbf{\bm{x}}_{i}\in\mathcal{X}$ is the nonsensitive feature of the $i$-th example, $y_{i}$ is its label, and $a_{i}$ is its sensitive feature. In a completely observed scenario, both $y_{i}$ and $a_{i}$ are 00 or $1$, and we will later extend it to missing values. As a shorthand, let $\mathbf{\bm{a}}=(a_{1},\ldots,a_{n})^{\top}$ and $\mathbf{\bm{y}}=(y_{1},\ldots,y_{n})^{\top}$. We first review some standard group fairness metrics:
Demographic parity (Dwork et al., 2012):
$P(\hat{Y}=1|A=0)=P(\hat{Y}=1|A=1),$ (1)
Equalized odds (Hardt et al., 2016): for $y\!\in\!\{0,\!1\}$
$\begin{split}P(\hat{Y}=1|A=0,Y=y)=\!P(\hat{Y}=1|A=1,Y=y),\!\end{split}\!$ (2)
Equal opportunity (Hardt et al., 2016):
$P(\hat{Y}=1|A=0,Y=1)=\!P(\hat{Y}=1|A=1,Y=1).\!\!$ (3)
In general, enforcing perfect fairness can be too restrictive, and approximate fairness can be considered via fairness measures. For example, the mean difference compares their difference (Calders and Verwer, 2010)
$\text{MD}(f)=P(\hat{Y}=1|A=0)-P(\hat{Y}=1|A=1).$ (4)
and the disparate impact (DI) factor computes their ratio (Feldman et al., 2015). Similar measures can be defined for equal opportunity and equalized odds. For continuous variables, $\chi^{2}$-divergence measures the independence as stipulated by separation and independence (Mary et al., 2019).
Given a dataset $\mathcal{D}_{tr}$ with $n$ examples, we can empirically estimate both sides of (1) by, e.g., for $a\in\{0,1\}$
$P(\hat{Y}=1|A=a)$ $\approx\text{mean}(\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=a\})\!\!\!\!\!$ (5)
$=\sum\nolimits_{i:a_{i}=a}P_{f}(\hat{Y}=1|x_{i})\Big{/}\sum\nolimits_{i:a_{i}=a}\!1.$
Here, for a finite set $S$, we denote the mean of its elements as $\text{mean}\,(S)$. The above estimator is straightforward from Eq 1 and 7 of Menon and Williamson (2018) in the setting of randomized classifier. It is asymptotically unbiased, consistent, and differentiable in $f$; see the proof in Appendix E. Similarly, we can estimate $P(\hat{Y}=1|A=a,Y=y)\approx\text{mean}\,(S),{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=a,y_{i}=y\}.$ Other estimators can also be adopted, such as kernel density estimation (Cho et al., 2020). Applying these estimates to MD, DI, $\chi^{2}$-divergence, or any other metric, we obtain a fairness risk $\mathcal{F}(P_{f},\mathbf{\bm{y}},\mathbf{\bm{a}})$. We keep its expression general, and some concrete examples are given in (18) to (20) in Appendix A.1. An example based on (5) is
$\mathcal{F}(P_{f},\mathbf{\bm{y}},\mathbf{\bm{a}})$ $:=|\text{mean}(\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=1\})$
$\qquad-\text{mean}(\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=0\})|.$
Classification risk.
We denote the conventional classification risk as $\mathcal{R}(P_{f},\mathbf{\bm{y}})$. For example, the standard cross-entropy loss yields $-\frac{1}{n}\sum_{i}\log P_{f}(\hat{Y}=y_{i}|X=x_{i})$. Combined with fairness risk, we can next find the optimal $f$ by minimizing their sum weighted by $\lambda>0$:
$\mathcal{F}(P_{f},\mathbf{\bm{y}},\mathbf{\bm{a}})+\lambda\cdot\mathcal{R}(P_{f},\mathbf{\bm{y}}).$ (6)
In Williamson and Menon (2019), a more general framework of fairness risk is constructed, incorporating the loss $\ell$ into the definition of $\mathcal{F}$ itself (Lahoti et al., 2020). Our method developed below can be applied directly.
4 Fairness with Group-Conditionally Unavailable Demographics
The regularized objective (6) requires the demographic features which can often become unavailable due to a respondent’s preference, privacy, and legal reasons. As a key contribution of this work, we further address group-conditionally unavailable demographics, where demographics can get unavailable in a non-uniform fashion, depending on the specific group. This matches reality. For example, people above some age may be more reluctant to reveal it when applying entry-level jobs.
We show that this new setting can be approached by extending the objective (6). To this end, we introduce a new random variable $\tilde{A}$, which is the observation of the true latent demographic $A$. For a principled treatment, we treat $A$ as never observed, while $\tilde{A}$ can be equal to $A$, or take a distorted value, or be unavailable (denoted as $\emptyset$). [^2] We intentionally term it “unavailable” instead of “missing”, because missing means unobserved in statistics, while we take unavailable (i.e., $\emptyset$) as an observed outcome. In other words, $\tilde{A}$ is always observed, including taking the value of $\emptyset$. As Section 5 shows, this novel model offers substantial flexibility, even if the demographic contains multiple features and a different subset of them is unavailable for different individuals.
We model the group-conditional unavailability with a learnable noising probability $P(\tilde{A}|A)$ (Yao et al., 2023). As a special case, one may assert that demographics cannot be misrepresented, i.e., $P(\tilde{A}|A)=0$ for all $\tilde{A}\notin\{A,\emptyset\}$. In order to address isometry (e.g., totally swapping the concept of male and female), we impose some prior on $P(\tilde{A}|A)$. For example, $P(\tilde{A}|A)$ must be low for $\tilde{A}\notin\{A,\emptyset\}$. This can be enforced by a Dirichlet prior, e.g., $(P(\text{male}|\text{male}),P(\text{female}|\text{male}),P(\emptyset|\text{male}))\sim\text{Dir}(0.5,0.1,0.4)$. Similarly, we can define $P(\tilde{Y}|Y)$ for class-conditionally unavailable label, allowing $\tilde{Y}=\emptyset$.
In the sequel, we will assume there is an SSL algorithm that predicts $Y$ and $A$ with probabilistic models $P_{f}(Y|X,\tilde{Y})$ and $q_{\phi}(A|X,\tilde{A})$, respectively. For example, the encoder of a VAE, which will be detailed in Section 5 as (5.1). Both $A$ and $Y$ can be multi-class. It is now natural to extend the fairness risk by taking the expectation of these missing values:
$\mathcal{E}_{\text{vanilla}}(P_{f},q_{\phi})$ (7)
$:=\mathop{\mathbb{E}}\limits_{Y_{i}\sim P_{f}(Y|x_{i},{\tilde{y}}_{i})}\mathop{\mathbb{E}}\limits_{A_{i}\sim q_{\phi}(A|x_{i},{\tilde{a}}_{i})}\mathcal{F}(P_{f},Y_{1:n},A_{1:n}).$
In the case of no misrepresentation, the expectation of $A_{i}$ can be replaced by $A_{i}={\tilde{a}}_{i}$ if ${\tilde{a}}_{i}\neq\emptyset$. Likewise for $Y_{i}$. Adding this vanilla semi-supervised regularizer to an existing learning objective of $P_{f}$ and $q_{\phi}(A|X,\tilde{A})$ such as VAE gives a straightforward promotion of fairness, and we will also discuss its optimization in Section 4.4. However, despite its clear motivation, we now demonstrate that it is indeed plagued with conceptual and practical issues, which we will address next.
4.1 Rationalizing semi-supervised fairness risk
It is important to note that the onus of fairness is only supposed to be on the classifier $P_{f}$, while $q_{\phi}(A|X,\tilde{A})$—which infers the demographics to define the fairness metric itself—is supposed to be recused from fairness risk minimization. Reducing the risk by manipulating an individual’s gender is not reasonable. This issue does not exist in a fully supervised setting, but complicates the regularizer when $q_{\phi}(A|X,\tilde{A})$ is jointly optimized with the classifier $P_{f}$. For example, although the risk can be reduced by making $A$ independent of $X$, this should not be intentionally pursued unless the underlying data distribution supports it. The posterior $q_{\phi}(A|X,\tilde{A})$ is only supposed to accurately estimate the demographics, while the task of fairness should be left to the classifier $P_{f}$.
A natural workaround is a two-step approach: first train a semi-supervised learner for the distribution $A|X,\tilde{A}$, and denote it as $r(A|X,\tilde{A})$. Then in (7), we replace $A_{i}\sim q_{\phi}(A|x_{i},{\tilde{a}}_{i})$ with $A_{i}\sim r(A|x_{i},{\tilde{a}}_{i})$. Although this resolves the above issue, it decouples the learning of $A|X,\tilde{A}$ (first step) and $Y|X,\tilde{Y}$ (second step). This is not ideal because, as commonly recognized, the inference of $A$ and $Y$ can benefit from shared backbone feature extractors on $X$ (e.g., ResNet for images), followed by finetuning target-specific heads.
However, sharing backbones will allow $q_{\phi}(A|X,\tilde{A})$ to be influenced by the learning of $P_{f}(Y|X,\tilde{Y})$, conflicting with the aforementioned recusal principle. In light of this difficulty, we resort to the commonly used stop-gradient technique that is available in PyTorch and TensorFlow. While they still share backbones, the derivative of $\mathcal{E}_{\text{vanilla}}$ in (7) with respect to $q_{\phi}(A|X,\tilde{A})$ is stopped (set to 0) from backpropagation, i.e., treating $q_{\phi}(A|X,\tilde{A})$ as a constant. It also much simplifies the training process.
4.2 Imputation of unavailable training labels
A similar issue is also present in the expectation $Y_{i}\sim P_{f}(Y|x_{i},{\tilde{y}}_{i})$ in (7). The ground truth label $Y_{1:n}$ is given in the supervised setting, but its missing values should not be inferred to account for fairness, because labels are part of the fairness definition itself, except for demographic parity. We could resort to the stop-gradient technique again to withhold the backpropagation through $Y_{i}\sim P_{f}(Y|x_{i},{\tilde{y}}_{i})$. However, different from unavailable demographics, the situation here is more intricate because $P_{f}$ also serves as the first argument of $\mathcal{F}(P_{f},Y_{1:n},A_{1:n})$, where it indeed enforces fair classification. This conflicts with the fairness-oblivious requirement for imputing $Y_{i}$, which is also based on $P_{f}$. In other words, we cannot enforce $P_{f}$ to both respect and disregard fairness at the same time.
To address this issue, we will train a separate semi-supervised classifier to impute $Y_{i}$ without fairness concerns, and it does leverage the backbone features mentioned above. Suppose $\mathbf{\bm{h}}_{i}$ is the current feature representation for $x_{i}$, e.g., the third last layer of $f$. Assuming there is no misrepresentation in ${\tilde{y}}_{i}$, we can clamp $Y|x_{i}$ to $P_{f}$ for those examples with observed label (${\tilde{y}}_{i}\neq\emptyset$). Then we infer a class probability ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{g}}(Y|x_{i})$ for the rest examples by using $\mathbf{\bm{h}}_{i}$ and any SSL algorithm such as graph-based Gaussian field (Zhu et al., 2003). This amounts to a fairness risk as
$\boxed{\mathcal{E}(P_{f},q_{\phi})}$ (8)
$:=\!\!\mathop{\mathbb{E}}\limits_{\{Y_{i}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{g}}(Y|x_{i}):{\tilde{y}}_{i}=\emptyset\}}\mathop{\mathbb{E}}\limits_{A_{i}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q_{\phi}}(A|x_{i},{\tilde{a}}_{i})}\!\!\mathcal{F}(P_{f},Y_{1:n},A_{1:n}).$
Note we stop the gradient on $P_{g}$ and $q_{\phi}$ here, and $P_{g}$ is used instead of $P_{f}$. Appendix A.6 extends this risk from randomized classifier to class-probability estimator.
(a) Decoder
(b) Encoder
Figure 1: SS-VAE decoder and encoder with unavailable demographic/label conditioned on group/class
4.3 Efficient evaluation of fairness risk $\mathcal{E}$
A major challenge in risk-based methods is the cost of computing the expectations in $\mathcal{E}$, along with its derivatives. In the sequel, we will resort to a Monte Carlo based method, such that for $n$ training examples, an $\epsilon$ accurate approximation with confidence $1-\delta$ can be found with $O(\frac{c_{n}}{\epsilon^{2}}\log\frac{1}{\delta})$ computation, where $c_{n}$ is the cost of evaluating $\mathcal{F}$ and it is $O(n)$ in our considered cases. Although some customized algorithms can be developed by exploiting the specific structures in $\mathcal{F}$ such as demographic parity, we prefer a more general approach. Despite the inevitable inexactness in the result, we prove tight concentration bounds that turn out sufficiently accurate in practice.
Our method simply draws $N$ number of iid samples from $A_{i}\sim q_{\phi}(A|x_{i},{\tilde{a}}_{i})$ and $Y_{i}\sim P_{g}(Y|x_{i})$. For $s=1,\ldots,N$, let $Z_{s}:=\{a^{(s)}_{i},y^{(s)}_{i}\}_{i=1}^{n}$. We estimate $\mathcal{E}$ by
^$\hat{\mathcal{E}}_{n}(Z_{1},\ldots,Z_{N}):={\textstyle\frac{1}{N}}\sum\nolimits_{s=1}^{N}\mathcal{F}(P_{f},Z_{s}).$ (9)
We prove the following sample complexity for this estimator by leveraging McDiarmid’s inequality.
Theorem 1.
Suppose $\mathcal{F}\in[0,C]$ where $C>0$ is a constant. Then for all $\epsilon>0$,
[…]
As a result, to guarantee an estimation error of $\epsilon$ with confidence $1-\delta$, it suffices to draw $N=\frac{C^{2}}{2\epsilon^{2}}\log\frac{1}{\delta}$ samples. The proof is available in Appendix A.2.
As ^$\hat{\mathcal{E}}_{n}(Z_{1},\ldots,Z_{N})$ costs $O(nN)$ to compute, the total cost is $\frac{nC^{2}}{2\epsilon^{2}}\log\frac{2}{\delta}$. When $C=1$, $\epsilon=0.01$ and $n=10^{4}$, this is order of $10^{8}$, which is quite affordable. In experiments, a sample size as low as $N=100$ was sufficient for our method to outperform the state of the art; see Section 6. Most risks we consider satisfy the boundedness assumption, although exceptions exist such as disparate impact.
4.4 Differentiation of vanilla fairness risk
Since we stop the gradient with respect to $q_{\phi}(A|x_{i},{\tilde{a}}_{i})$ and $P_{g}(Y|x_{i})$, the whole regularizer $\mathcal{E}(P_{f},q_{\phi})$ in (8) can be easily differentiated with respect to $f$.
In practice, we also would like to conduct an ablation study by comparing with $\mathcal{E}_{\text{vanilla}}$, where challenges arise from differentiation with respect to $q_{\phi}(A|x_{i},{\tilde{a}}_{i})$ and $y_{i}\sim P_{f}(Y|x_{i},{\tilde{y}}_{i})$. We resolve this issue by using the straight-through Gumbel-Softmax method (Jang et al., 2017; Maddison et al., 2017). A self-contained description is given in Appendix A.3.
5 Integrating Fairness Risk with SSL
We next illustrate how the fairness risk can be integrated with SSL. As an example, we recap the semi-supervised VAE (SS-VAE, Kingma et al., 2014) with missing $a$ and $y$ values. [^3] It is customary in statistics to denote random variables by capital letters. However, the VAE literature generally uses lowercase letters. We thus switch to their custom.
5.1 Encoders and decoders
SS-VAE employs a decoder/generation process and an encoder/inference process parameterized by $\theta$ and $\phi$ respectively. In the decoder whose graphical model is shown in Figure 1(a), the observation $\mathbf{\bm{x}}$ is generated conditioned on the latent variable $\mathbf{\bm{z}}$, latent class $y$ and latent demographics $a$:
$p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}|a,y,\mathbf{\bm{z}})$ $=\mathcal{N}(\mathbf{\bm{x}}|g_{\theta}(a,y,\mathbf{\bm{z}}))\cdot P_{\theta}({\tilde{a}}|a)\cdot P_{\theta}({\tilde{y}}|y),$
$p(y)$ $=\text{Cat}(y|\pi_{y}),\quad p(\pi_{y})=\text{SymDir}(\gamma_{y}),$
$p(a)$ $=\text{Cat}(a|\pi_{a}),\quad p(\pi_{a})=\text{SymDir}(\gamma_{a}),$
$p(\mathbf{\bm{z}})$ $=\mathcal{N}(\mathbf{\bm{z}}|0,I).$ (10)
Here $\text{Cat}(y|\pi_{y})$ is the multinoulli prior of the class variable. SymDir is the symmetric Dirichlet with hyperparameter $\gamma_{y}$. The prior of $a$ is analogously defined.
The inference process finds conditionally independent representations $\mathbf{\bm{z}}$, $a$, and $y$ under a given $\mathbf{\bm{x}}$, ${\tilde{a}}$, and ${\tilde{y}}$. Accordingly, the approximate posterior/encoder $q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ factors as in Figure 1(b):
$q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ $=q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})\cdot q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})\cdot q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})\qquad$
$\text{where}\ \ \quad q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})$ $=\mathcal{N}(\mathbf{\bm{z}}|\mu_{\phi}(\mathbf{\bm{x}}),\sigma^{2}_{\phi}(\mathbf{\bm{x}})),\quad$
$q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})$ $\propto q_{\phi}(y|\mathbf{\bm{x}})\cdot P_{\theta}({\tilde{y}}|y)$
$=\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}}))\cdot P_{\theta}({\tilde{y}}|y),$
$q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})$ $\propto q_{\phi}(a|\mathbf{\bm{x}})\cdot P_{\theta}({\tilde{a}}|a)$
$=\text{Cat}(a|h_{\phi}(\mathbf{\bm{x}}))\cdot P_{\theta}({\tilde{a}}|a).$ (11)
Here $\mu_{\phi},\sigma_{\phi},g_{\phi},h_{\phi}$ are all defined as neural networks. Compared with Figure 1(a), here we only reversed the arrows connected with $\mathbf{\bm{x}}$, following a standard assumption in VAE that, in posterior, $y$, $z$, $a$ are independent given $\mathbf{\bm{x}}$. We maximally preserved the other arrows, namely $y\to{\tilde{y}}$ and $a\to{\tilde{a}}$. The chain of $\mathbf{\bm{x}}\to a\to{\tilde{a}}$ can be interpreted as first probabilistically determining $a$ based on $\mathbf{\bm{x}}$, and then adding noise to it producing ${\tilde{a}}$ (including $\emptyset$). This chain structure allows us to derive $q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})$ in (5.1) (and analogously $q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})$) as:
$p(a|\mathbf{\bm{x}},{\tilde{a}})$ $\propto p(a)p(\mathbf{\bm{x}},{\tilde{a}}|a)=p(a)p(\mathbf{\bm{x}}|a)p({\tilde{a}}|a)$
$=p(\mathbf{\bm{x}})p(a|\mathbf{\bm{x}})p({\tilde{a}}|a)\propto p(a|\mathbf{\bm{x}})p({\tilde{a}}|a).$
In training, the fairness risks in (7) and (8) require $P_{f}(y|\mathbf{\bm{x}},{\tilde{y}})$, which can be served by $q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})$.
5.2 Instilling Fairness to SS-VAE
The evidence lower bound (ELBO) of $\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ can be derived in a standard fashion. Denoting it as $\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$, we can extend it as
$\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\geq\!\!\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{z}},a,y)\sim q_{\phi}(\cdot|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})}[\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}|a,y,\mathbf{\bm{z}})\!$
$+\log(p(y)p(a)p(\mathbf{\bm{z}}))-\log q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})]$
$=\mathop{\mathbb{E}}\limits_{\mathbf{\bm{z}}\sim q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})}\log p_{\theta}(x|a,y,\mathbf{\bm{z}})+\mathop{\mathbb{E}}\limits_{y\sim q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})}\log P_{\theta}({\tilde{y}}|y)$
$+\mathop{\mathbb{E}}\limits_{a\sim q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})}\log P_{\theta}({\tilde{a}}|a)-\text{KL}[q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})||p(\mathbf{\bm{z}})]$
$-\text{KL}[q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})||p(y)]-\text{KL}[q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})||p(a)]$
$=:\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}).$
Followed by ELBO, we extend the SS-VAE objective as follows, to be minimized over $\theta$ and $\phi$:
$\mathcal{L}(\theta,\phi)$ $=\Omega(P_{\theta}({\tilde{a}}|a),P_{\theta}({\tilde{y}}|y))$
$-\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\sim\mathcal{D}_{tr}}\!\!\Big{[}\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ (12)
$+[\![{\tilde{a}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{a}\text{Cat}(a|h_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{a}}|a)$
$+[\![{\tilde{y}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{y}\!\!\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{y}}|y)\Big{]}.$
Here, the regularizer $\Omega$ enforces the Dirichlet prior. Suppose ${\tilde{y}}\neq\emptyset$. If the provided ${\tilde{y}}$ always truthfully represents $y$, then $P({\tilde{y}}|y)=[\![{\tilde{y}}=y]\!]$. As a result, the last line in (12) equals $\text{Cat}({\tilde{y}}|g_{\phi}(\mathbf{\bm{x}}))$, the standard supervised loss in SS-VAE.
Casting SS-VAE into the fairness risk framework, the role of $f$ is played by $q_{\phi}(y|\mathbf{\bm{x}})$ and we only need to augment the objective (12) into
$\boxed{\mathcal{L}(\theta,\phi)+\lambda\cdot\mathcal{E}(q_{\phi}(y|\mathbf{\bm{x}}),q_{\phi}(a|\mathbf{\bm{x}})),}$ (13)
where $\lambda>0$ is a tradeoff hyperparameter. We will henceforth refer to this model as Fair-SS-VAE.
Classifying test data.
We simply predict by
$P_{f}(y|\mathbf{\bm{x}}):=q_{\phi}(y|\mathbf{\bm{x}})=\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}})),$ (14)
with no access to ${\tilde{a}}$ (Lipton et al., 2018). A more principled approach is to minimize the risk of fairness and classification, which is relegated to Appendix A.5.
Note our new VAE can readily extend $y$ and $a$ from binary values to multi-class and continuous, with little impact on Monte Carlo sampling in (9). So our Fair-SS-VAE is applicable wherever the underlying fairness risk is available; see Appendix A.1.
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 2: Pareto frontier of error versus DEO/DEOPP for Adult-Gender
6 Experimental Results
We now show empirically that with group-conditionally missing demographics and generally missing labels, our Fair-SS-VAE significantly outperforms various state-of-the-art semi-supervised fair classification methods in terms of various fairness metrics, while keeping a similar accuracy.
Datasets
We used two datasets to create three sets of experiments:
CelebA (Liu et al., 2015). CelebA contains face images annotated with 40 binary attributes. We sampled 45k images as our training and validation dataset, and 5k as test set. Following Jung et al. (2022), we set ”Attractive” as the target label, and ”Gender” as the sensitive group attribute.
UCI Adult (Becker and Kohavi, 1996). The Adult dataset is a tabular dataset where the target label is whether the income exceeds \$50K per year given a person’s attributes. We conducted two experiments with gender and race serving the sensitive attribute. The same pre-processing routine as in Bellamy et al. (2018) was adopted.
To simulate the missing data, we randomly masked 25% labels, and masked the demographics as:
$P(\tilde{A}\!=\!\emptyset|A\!=\!1)=\alpha,\ \ P(\tilde{A}\!=\!1|A\!=\!1)=1\!-\!\alpha,$
$P(\tilde{A}\!=\!\emptyset|A\!=\!0)=\beta,\ \ P(\tilde{A}\!=\!0|A\!=\!0)=1\!-\!\beta.$
We considered three levels of missing demographics: sparse ($\alpha=0.4,\beta=0.8$), medium ($\alpha=0.2,\beta=0.4$), and dense ($\alpha=0.1,\beta=0.2$). Fair-SS-VAE can infer $\alpha$ and $\beta$ from the data. Each dataset was first randomly partitioned into training, validation, and testing. Masking was applied to the first two, and no group information is used in testing.
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 3: Pareto frontier of error versus DEO/DEOPP for CelebA
Baseline fair classifiers
We compared with three types of methods. MMD-based Fair Distillation (MFD, Jung et al., 2021) encourages fairness through feature distillation, but it requires observed labels and demographics. So we imputed $y$ and $a$ with three heuristics:
  • (gt y, gt a): feed the ground-truth $y$ and $a$. It forms an unfair comparison with our method but sheds light on the best possible performance of the method;
  • (gt y, rand a): feed the ground-truth $y$ and impute $a$ uniformly at random to eliminate the correlation between labels and groups;
  • (pred y, pred a): train a classifier to predict the missing $y$ and $a$;
As plotting three curves makes the figures overly crowded, we defer the full plots to Appendix C, and only show in the main paper the results without involving ground-truth imputation (the last variant).
The second baseline is FairRF (Zhao et al., 2022), which uses proxy features. For Adult, they used age, relation and marital status as the proxy features. We found empirically that using age alone gave even better results. So we just used age as the proxy feature. We did not apply FairRF to CelebA because proxy features are difficult to identify for images.
The third baseline is FairDA (Liang et al., 2023). It divides all examples into source and target domains based on the availability of sensitive feature, and estimates that for the target domain by domain adaptation.
Implementation details of Fair-SS-VAE
We employed M2-VAE (Kingma et al., 2014) for Fair-SS-VAE, with WideResNet-28-2 used as the feature extractor for CelebA, and three fully connected layers for Adult. We kept the sample size to 100 in the Monte-Carlo evaluation of $\mathcal{E}$ in (9).
Fairness performance metrics
We measured the fairness of the predictions by difference of equalized odds (DEO) and difference of equal opportunity (DEOPP). On a test set, the prediction $P(\hat{Y}=1|x_{i})$ is valued 0 or 1, simply rounded from $q_{\phi}(y|x)$. Surprisingly, it provides a very strong initialization for minimizing the expected fairness risk in (A.5). In fact, it directly hits a local optimal, leaving no improvement possible from coordinate descent. This suggests that the fairness regularization in Fair-SS-VAE allows VAE to learn a posterior model that already accounts for the desired fairness, dispensing with the need of risk minimization at test time.
Model selection
Since we will plot the Pareto frontier of classification error and fairness metrics, ideally we only need a training and test set, and to enumerate all the hyperparameter values, based on which the Pareto frontier can be plotted. However, doing so would be too expensive in computation. Therefore, we used 10% of the data to tune all the other hyper-parameters (network architecture, optimization, etc), and varied the tradeoff weight $\lambda$ in (13) to roll out the frontier.
a
6.1 Main results
Figures 2 and 3 show respectively the Pareto frontier of classification error and fairness metrics for Adult-Gender (with gender being the sensitive feature) and CelebA. Due to space limitation, we defer the result of Adult-Race to Figure A.2 in Appendix C. As the missing demographics decrease from left to right (sparse to dense), the Pareto Frontier tends to shift from the upper right corner to the bottom left corner, suggesting that all methods show improvement in the fairness metric (both DEO and DEOPP) while maintaining the same level of accuracy.
Fair-SS-VAE excels in producing a Pareto Frontier significantly closer to the origin across diverse datasets and varying levels of sparsity, indicating its superiority in balancing the tradeoff between accuracy and the fairness metric, unless the ground-truth $y$ and $a$ are used, which would make the comparison unfair to Fair-SS-VAE (see Appendix C). Overall, FairDA is the second most effective.
Recovery of group-conditional missing probability
Table 1 presents the value of $\alpha$ and $\beta$ that Fair-SS-VAE finds for the three datasets and three sparsity levels. Such a recovery is difficult because, once masked, the data does not carry the ground-truth of group feature, precluding counting for rate estimation. Although the exact values are hard to recover, in many cases, they turn out close, and the overall trend correctly decreases from sparse to dense.
Table 1: Recovery rate of sensitive attribute by Fair-SS-VAE. Ad-G: Adult-Gender, Ad-R: Adult-Race, CelA: CelebA
Sparse Medium Dense
[…]
Ad-G .43$\pm$.16 .86$\pm$.17 .32$\pm$.06 .35$\pm$.06 .11$\pm$.07 .28$\pm$.04
Ad-R .39$\pm$.23 .82$\pm$.15 .30$\pm$.04 .42$\pm$.13 .17$\pm$.01 .25$\pm$.12
CelA .34$\pm$.16 .67$\pm$.02 .08$\pm$.02 .44$\pm$.13 .16$\pm$.04 .29$\pm$.04
Ablation study of stopping gradient
To investigate the significance of stop gradient, we set up an ablation experiment that directly uses $\mathcal{E}_{\text{vanilla}}$. The corresponding results are labeled “Ablation” in Figure 2, 3 and A.2. Under most circumstances, it performs worse than Fair-SS-VAE, with occasional advantages when applied to datasets with dense sparsity levels.
Conclusion, limitation, broader impact, and future work
We proposed a new fair classifier that addresses group-conditionally missing demographics. Promising empirical performance is shown. As a limitation, we did not incorporate proxy features when they are available. For future work, we will enable it by conditioning the VAEs on the proxy features. We will also model the bias between different sensitive groups, e.g., people of some race are less reluctant to reveal their gender.
Supplementary Materials
Appendix A Conceptual Details
In this appendix section, we fill in more technical details and proofs from the main paper.
A.1 Multi-class and Continuous Valued Fairness Risks
Fairness metrics can be extended to multi-class labels and sensitive features in a number of different ways (Rouzot et al., 2022; Xian and Zhao, 2023). Here we recap the disparity of true positive rates (TPRs). Let the label space be $[n_{y}]:=\{1,\ldots,n_{y}\}$, and the sensitive feature space be $[n_{a}]$. Denote
$\text{TPR}(y,a):=P(\hat{Y}=y|Y=y,A=a),\forall\ y\in[n_{y}],\ a\in[n_{a}].$
Let $S_{y}$ be the set of labels of interest, e.g., $S_{y}=[n_{y}]$. Then we can define the fairness metric as
$\Delta_{\text{TPR}}:=\max_{a,a^{\prime}\in[n_{a}]}\max_{y,y^{\prime}\in S_{y}}\left|P(\hat{Y}=y|Y=y,A=a)-P(\hat{Y}=y^{\prime}|Y=y^{\prime},A=a^{\prime})\right|.$ (15)
When $\Delta_{\text{TPR}}=0$ and the classes are binary, this fairness notion recovers equalized odds with $S_{y}=[n_{y}]$, equal opportunity with $S_{y}=\{1\}$, and predictive equality with $S_{y}=\{0\}$ (Chouldechova, 2017).
Suppose there is a classifier $f$ that assigns the probability $P_{f}(\hat{Y}=y|x)$. Then we can use the data to estimate
$P(\hat{Y}=y|Y=y,A=a)\approx\text{mean}\,S,{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=y|x_{i}):a_{i}=a,y_{i}=y\}.$
Plugging it into $\Delta_{\text{TPR}}$ would directly provide a differentiable fairness risk $\mathcal{F}$ for regularization.
Demographic parity:
we need $\hat{Y}$ to be independent of $A$, and the degree of dependence can be measured by any existing applicable metric such as mutual information or
$\max_{a,a^{\prime}\in[n_{a}]}\max_{y\in S_{y}}\left|P(\hat{Y}=y|A=a)-P(\hat{Y}=y|A=a^{\prime})\right|$ (16)
or $\max_{a\in[n_{a}]}\max_{y\in S_{y}}\left|P(\hat{Y}=y|A=a)-P(\hat{Y}=y)\right|$ (17)
Again, we can estimate the relevant probabilities from data by
$P(\hat{Y}=y|A=a)\approx\text{mean}\,S,{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=y|x_{i}):a_{i}=a\}$
$P(\hat{Y}=y)\approx\text{mean}\,S,{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=y|x_{i})\}.$
Concrete fairness risk expressions
For ease of reference, we explicitly write out the expressions.
For equalized odds
$\mathcal{F}_{\text{DEO}}$ $=\max_{y\in\{0,1\}}|\text{mean}\{P_{f}(\hat{Y}=y|x_{i}):a_{i}=0,y_{i}=y\}-\text{mean}\{P_{f}(\hat{Y}=y|x_{i}):a_{i}=1,y_{i}=y\}|$
$=\max_{y\in\{0,1\}}|\text{mean}\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=0,y_{i}=y\}-\text{mean}\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=1,y_{i}=y\}|.$ (18)
For equal opportunity:
$\mathcal{F}_{\text{DEOPP}}=|\text{mean}\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=0,y_{i}=1\}-\text{mean}\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=1,y_{i}=1\}|.$ (19)
For demographic parity:
$\mathcal{F}_{\text{DDP}}$ $=\max_{y\in\{0,1\}}|\text{mean}\{P_{f}(\hat{Y}=y|x_{i}):a_{i}=0\}-\text{mean}\{P_{f}(\hat{Y}=y|x_{i}):a_{i}=1\}|$
$=|\text{mean}\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=0\}-\text{mean}\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=1\}|.$ (20)
Continuously valued label and sensitive features
When the label and sensitive features are continuous, the above methods cease to be applicable. We will resort to variational estimators such as InfoNCE, and they require paired sampled of $\hat{Y}$ and $A$. However, we can do better than that because the classifier $f$ provides a distribution of $\hat{Y}$ instead of a single sample. For example, in mutual information neural estimation (MINE), we can estimate the KL-divergence between $P_{A,\hat{Y}}$ and $P_{A}P_{\hat{Y}}$ by
[…]
Then we can estimate both terms by
[…]
If the function space of $t$ is simple enough, we can directly evaluate the two expectations in a closed form. Otherwise, we can sample $\hat{Y}$ from $P_{f}(\hat{Y}|x_{j})$ and apply Gumbel-softmax for differentiation.
A.2 Proof of Theorem 1
Proof.
First we evaluate the bounded difference: replace $Z_{s}$ with $Z^{\prime}_{s}$ while keeping all other $Z_{s^{\prime}}$ intact. Here $Z^{\prime}_{s}$ is any admissible assignment of $a^{(s)}_{i}$ and $y^{(s)}_{i}$. Since $\mathcal{F}$ takes value in $[0,1]$, it is trivial that |^$|\hat{\mathcal{E}}_{n}(Z_{1},\ldots,Z_{N})-\hat{\mathcal{E}}_{n}(Z_{1},\ldots,Z_{s-1},Z^{\prime}_{s},Z_{s+1},\ldots,Z_{N})|=\frac{1}{N}|\mathcal{F}(P_{f},Z_{s^{\prime}})-\mathcal{F}(P_{f},Z_{s})|\leq C/N$. Further noting that the expectation of ^$\hat{\mathcal{E}}_{n}(Z_{1},\ldots,Z_{N})$ over $(Z_{1},\ldots,Z_{N})$ is $\mathcal{E}$, the McDiarmid’s inequality immediately implies the theorem. ∎
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 4: Pareto frontier of error versus DEO/DEOPP for Adult-Race
A.3 Straight-through Gumbel-Softmax for vanilla fairness risk
Suppose a Bernoulli variable $y_{i}$ has $P(y_{i}=1)=q_{i}$. Then the vanilla Gumbel-Softmax method first draws i.i.d. samples from a Gumbel$(0,1)$ distribution:
$\{\alpha^{(s)}_{i},\beta^{(s)}_{i}:i=1,\ldots,n,\text{ and }\ s=1,\ldots,N\}.$ (22)
Then a sample of $y^{(s)}_{i}$ is constructed by
$y^{(s)}_{i}=\frac{\exp(\frac{\log q_{i}+\alpha^{(s)}_{i}}{T})}{\exp(\frac{\log q_{i}+\alpha^{(s)}_{i}}{T})+\exp(\frac{\log(1-q_{i})+\beta^{(s)}_{i}}{T})}\ ,$ (23)
where $T>0$ is a temperature parameter. A larger value of $T$ leads to a more smooth and uniform sample, i.e., $y^{(s)}_{i}$ will be closer to 0.5. Clearly $y^{(s)}_{i}$ is not discrete, but for reasonably small $T$, it will be close to 0 or 1. To patch up the non-integrality of $y^{(s)}_{i}$, the straight-through Gumbel-Softmax gradient estimator only uses the differentiable variable in the backward gradient propagation, while the forward pass still uses categorical variables (i.e., turning $y^{(s)}_{i}$ into 1 if it is above 0.5, and 0 otherwise). The same approach can be applied to sample $a_{i}$.
A.4 Soften the step function to enable differentiation
The step function [[>0]][\![$\cdot$>0]\!][ [ ⋅ > 0 ] ] offers no useful derivative. Noting that taking a threshold of a probability $f(X)$ is equivalent to thresholding its logit $\sigma^{-1}(f(X))$, we improve numerical performance by setting $P_{f}(Y=1|X)=\sigma(\frac{\sigma^{-1}(f(X))-\tau}{T})$, where the temperature $T$ controls the steepness of the approximation. Finally, group-specific threshold can be enabled by replacing $\tau$ with $\tau_{1}a_{i}+\tau_{0}(1-a_{i})$. Given the values of $a_{i}$—either from the training data or from the $N$ samples—it is straightforward to optimize this objective.
A.5 Labeling test data by minimizing fairness and classification risks
With the learned $\phi$, the test data can be labeled by simply seeking the $\mathbf{\bm{y}}$ that maximizes the likelihood while also minimizing the expected fairness risk:
$\mathrm{arg}\mathop{\mathrm{min}}_{y_{i}\in\{0,1\}}\Bigl{\{}-\frac{1}{n_{\text{test}}}\sum_{i}\log q_{\phi}(y_{i}|\mathbf{\bm{x}}_{i})+\mathop{\mathbb{E}}\limits_{A_{i}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}q_{\phi}}(A|\mathbf{\bm{x}}_{i},{\tilde{a}}_{i})}\mathcal{F}(P_{q_{\phi}(y|\mathbf{\bm{x}})},\mathbf{\bm{y}},A_{1:n_{\text{test}}})\Bigr{\}}.$
When all demographics are observed, this can be solved by tuning group-specific thresholds on $q_{\phi}(y_{i}=1|\mathbf{\bm{x}}_{i})$. When test demographics are missing (possibly in their entirety), we can first compute the binary demographics via $\mathrm{arg}\mathop{\mathrm{max}}_{A}q_{\phi}(A|\mathbf{\bm{x}}_{i},{\tilde{a}}_{i})$, and then tune the group-specific threshold on $q_{\phi}(y_{i}=1|\mathbf{\bm{x}}_{i})$. Afterwards, we do a coordinate descent to finetune $\mathbf{\bm{y}}$ in the face of full expectation on $A_{i}$ for $a_{i}=\emptyset$ using Monte Carlo.
A.6 Extension to class-probability estimation
The above discussion has been intentionally kept general by considering a randomized classifier for $Y|X$. In practice, however, this is not quite realistic because one does not get to predict multiple times for a single test set. Therefore, we would like to specialize the framework to classification based on class-probability, i.e., $P_{f}(Y=1|x)=[\![f(x)-\tau>0]\!]$. It also provides the convenience of tuning the parameter $\tau$. Once $h$ is learned and fixed thereafter, different test scenarios may demand different trade-offs between risks in classification and fairness. For example, Jang et al. (2022) proposed adapting the thresholds for each demographic group. To enable differentiation, the [[]]delimited-[]delimited-[][\![$\cdot$]\!][ [ ⋅ ] ] operator can be softened as shown in Appendix A.
Appendix B Experiment Settings
We implemented our algorithm using Python 3.11.5 and PyTorch 2.1.1. The code runs on a server with Ubuntu 18.04, powered by an Intel(R) Core(TM) i9-9900X CPU @ 3.50GHz and 128 GB of RAM. For GPU resources, we have four GeForce RTX 2080 Ti cards, each with 11 GB of graphic memory, and the GPU driver version is 455.23.05.
Appendix C Additional Experimental Results
Figure A.2 shows the Pareto frontier of error versus DEO/DEOPP for Adult-Race.
We additionally present the Pareto frontier attained by all the baseline methods. Figures C to C are crowded, hence moved to the appendix. The curve of mfd_gty_gta remains intact across all sparsity levels for a given dataset, because it imputes the sensitive feature with its ground truth. Note that it is not fair to compare Fair-SS-VAE against mfd_gty_gta, because the latter has full access to the ground truth of both labels and sensitive features. Nonetheless, we still visualize its Pareto frontier because it serves as a rough upper bound for all the methods.
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 5: Pareto frontier of error versus DEO/DEOPP for Adult-Gender
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 6: Pareto frontier of error versus DEO/DEOPP for Adult-Race
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 7: Pareto frontier of error versus DEO/DEOPP for CelebA
Appendix D Source Code
The experiment code is available at https://tinyurl.com/4wndtfbb.
Appendix E Statistical Property of Estimator (5)
Theorem 2.
The following estimator of $\theta:=P(\hat{Y}=1|A=a)$ is consistent and asymptotically unbiased:
$\theta_{n}:=\sum\nolimits_{i:a_{i}=a}P(\hat{Y}=1|x_{i})\Big{/}\sum\nolimits_{i:a_{i}=a}1\,.$ (24)
However, it is in general not unbiased.
Proof.
To show consistency, note
[…]
Therefore,
[…]
To show asymptotic unbiasedness, note
$\mathbb{E}[|\theta_{n}-\theta|]$ $=\mathbb{E}[|\theta_{n}-\theta|\cdot[\![|\theta_{n}-\theta|<\epsilon]\!]]+\mathbb{E}[|\theta_{n}-\theta|\cdot[\![|\theta_{n}-\theta|\geq\epsilon]\!]]$ (28)
$\leq\epsilon+\mathbb{E}[|\theta_{n}-\theta|\cdot[\![|\theta_{n}-\theta|\geq\epsilon]\!]]$ (29)
(by Cauchy-Schwarz) $\leq\epsilon+\sqrt{\mathbb{E}[|\theta_{n}-\theta|^{2}]\Pr(|\theta_{n}-\theta|\geq\epsilon)}.$ (30)
Consistency has already established that $\Pr(|\theta_{n}-\theta|\geq\epsilon)\to 0$ as $n\to\infty$. Since $\theta_{n}$ is clearly bounded by 1, the right-hand side falls below $2\epsilon$ for sufficiently large $n$. ∎

### v3

@@
 The above estimator is straightforward from Eq 1 and 7 of Menon and Williamson (2018) in the setting of randomized classifier.
-It is asymptotically unbiased, consistent, and differentiable in $f$; see the proof in Appendix E.
+It is asymptotically unbiased, consistent, and differentiable in $f$.
+see the proof in Appendix E.
 Similarly, we can estimate $P(\hat{Y}=1|A=a,Y=y)\approx\text{mean}\,(S),{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=a,y_{i}=y\}.$ Other estimators can also be adopted, such as kernel density estimation (Cho et al., 2020).
@@
 However, sharing backbones will allow $q_{\phi}(A|X,\tilde{A})$ to be influenced by the learning of $P_{f}(Y|X,\tilde{Y})$, conflicting with the aforementioned recusal principle.
-In light of this difficulty, we resort to the commonly used stop-gradient technique that is available in PyTorch and TensorFlow.
+In light of this difficulty, we resort to the commonly used stop-gradient technique that is available in PyTorch (detach) and TensorFlow (tf.stop_gradient).
 While they still share backbones, the derivative of $\mathcal{E}_{\text{vanilla}}$ in (7) with respect to $q_{\phi}(A|X,\tilde{A})$ is stopped (set to 0) from backpropagation, i.e., treating $q_{\phi}(A|X,\tilde{A})$ as a constant.
 It also much simplifies the training process.
+The stop-gradient method is a technique that blocks the flow of gradient through a specific part of the computational graph during backpropagation, which is commonly used to decouple certain computations from parameter updates.
+We present the corresponding results in our ablation study, shown by the blue curve among all experiment figures.
 4.2 Imputation of unavailable training labels
@@
 The proof is available in Appendix A.2.
+This theorem is non-trivial because the fairness risk can depend on the predicted probabilities in a complex manner, creating nonlinear dependencies among all the training examples; see Appendix A.1.
+The analysis also reveals the limitations of the sampler, as certain risks, such as disparate impact, are unbounded.
 As ^$\hat{\mathcal{E}}_{n}(Z_{1},\ldots,Z_{N})$ costs $O(nN)$ to compute, the total cost is $\frac{nC^{2}}{2\epsilon^{2}}\log\frac{2}{\delta}$.
@@
 We thus switch to their custom.
+Why VAE?
+In terms of SSL, while some discriminative methods exist such as Zhu et al.
+(2003), the most prevalent and effective approaches are based on generative models, such as SS-VAE.
+Indeed, most conventional generative models for SSL are essentially VAEs, especially when they are trained with variational inference.
+Secondly, our approach to fairness risk relies on an encoder capable of estimating the posterior probability of missing labels or demographics.
+This rules out normalizing flows or generative adversarial networks unsuitable, as the latter’s discriminator does not serve as an encoder.
+Moreover, transformers are not considered since our formulation does not involve sequential data modeling.
+In contrast, VAE employs an encoder and provides the conditional independencies that allow us to more finely specify the constituent distributions as in (10) and (5.1).
+Diffusion models, as a more refined variant of VAE, can also be adopted.
+However, we intentionally keep the generative model simple, prioritizing fairness modeling, while future work can incorporate additional refinements to the VAE.
 5.1 Encoders and decoders
@@
 5.2 Instilling Fairness to SS-VAE
-The evidence lower bound (ELBO) of $\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ can be derived in a standard fashion.
-Denoting it as $\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$, we can extend it as
-$\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\geq\!\!\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{z}},a,y)\sim q_{\phi}(\cdot|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})}[\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}|a,y,\mathbf{\bm{z}})\!$
-$+\log(p(y)p(a)p(\mathbf{\bm{z}}))-\log q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})]$
-$=\mathop{\mathbb{E}}\limits_{\mathbf{\bm{z}}\sim q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})}\log p_{\theta}(x|a,y,\mathbf{\bm{z}})+\mathop{\mathbb{E}}\limits_{y\sim q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})}\log P_{\theta}({\tilde{y}}|y)$
-$+\mathop{\mathbb{E}}\limits_{a\sim q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})}\log P_{\theta}({\tilde{a}}|a)-\text{KL}[q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})||p(\mathbf{\bm{z}})]$
-$-\text{KL}[q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})||p(y)]-\text{KL}[q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})||p(a)]$
-$=:\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}).$
-Followed by ELBO, we extend the SS-VAE objective as follows, to be minimized over $\theta$ and $\phi$:
+The evidence lower bound (ELBO) of $\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ can be derived in a standard fashion, and we relegate the details to Appendix A.4.
+Denoting it as $\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$, we extend the SS-VAE objective as follows, to be minimized over $\theta$ and $\phi$:
 $\mathcal{L}(\theta,\phi)=\Omega(P_{\theta}({\tilde{a}}|a),P_{\theta}({\tilde{y}}|y))$
-$-\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\sim\mathcal{D}_{tr}}\!\!\Big{[}\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ (12)
-$+[\![{\tilde{a}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{a}\text{Cat}(a|h_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{a}}|a)$
-$+[\![{\tilde{y}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{y}\!\!\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{y}}|y)\Big{]}.$
+$\qquad\qquad-\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\sim\mathcal{D}_{tr}}\!\!\Big{[}\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ (12)
+$\qquad\qquad\qquad+[\![{\tilde{a}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{a}\text{Cat}(a|h_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{a}}|a)$
+$\qquad\qquad\qquad+[\![{\tilde{y}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{y}\!\!\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{y}}|y)\Big{]}.$
 Here, the regularizer $\Omega$ enforces the Dirichlet prior.
+For example, let $a=P(\text{male}|\text{male})$, $b=P(\text{female}|\text{male})$, $c=P(\emptyset|\text{male})$.
+Then a prior of $\text{Dir}(0.5,0.2,0.4)$ leads to a regularizer of $\log(a^{0.5-1}b^{0.2-1}c^{0.4-1})$.
+The regularization is beneficial when prior knowledge about the distributions is available, such as the likelihood that a male is more likely to be labeled as male rather than female.
+However our experiments did not make any such assumptions and thus did not incorporate this regularizer.
 Suppose ${\tilde{y}}\neq\emptyset$.
@@
 $\boxed{\mathcal{L}(\theta,\phi)+\lambda\cdot\mathcal{E}(q_{\phi}(y|\mathbf{\bm{x}}),q_{\phi}(a|\mathbf{\bm{x}})),}$ (13)
-where $\lambda>0$ is a tradeoff hyperparameter. We will henceforth refer to this model as Fair-SS-VAE.
+where $\lambda>0$ is a tradeoff hyperparameter and $\mathcal{E}$ takes the same form as in (8). We will henceforth refer to this model as Fair-SS-VAE.
 Classifying test data.
@@
 We sampled 45k images as our training and validation dataset, and 5k as test set.
-Following Jung et al.
-(2022), we set ”Attractive” as the target label, and ”Gender” as the sensitive group attribute.
+Following (Jung et al., 2022), we set ”Attractive” as the target label, and ”Gender” as the sensitive group attribute.
 UCI Adult (Becker and Kohavi, 1996).
@@
 (2018) was adopted.
+It is worth mentioning that conducting experiments where there is a significant correlation between target labels and groups is necessary.
+If the labels and groups are already uncorrelated, fairness modeling is likely less critical.
+For example, in CelebA, when considering a label/group pair with very low correlation, such as ’Eyeglasses’ and ’Male’, the model already achieves high accuracy with a DEO close to zero (ranging from 0.000 to 0.003) even without any fairness interventions.
+This suggests that fairness-aware models, including ours, have limited potential for improvement, making them less ideal for experiments.
 To simulate the missing data, we randomly masked 25% labels, and masked the demographics as:
@@
 Therefore, we used 10% of the data to tune all the other hyper-parameters (network architecture, optimization, etc), and varied the tradeoff weight $\lambda$ in (13) to roll out the frontier.
-a
+Precisely, we focused on varying the tradeoff weight between accuracy and fairness while keeping other hyperparameters fixed, such as batch size, learning rate, and the weights for KL divergence and reconstruction error in the VAE.
+Since these hyperparameters are already part of the standard SS-VAE, we relied on SS-VAE to select their values.
+After determining these values, we put back the fairness risk and further adjusted the selected hyperparameters, generating a new set of candidate values for model selection based on the validation set.
 6.1 Main results
@@
 Under most circumstances, it performs worse than Fair-SS-VAE, with occasional advantages when applied to datasets with dense sparsity levels.
+Additionally, we carried out experiments titled CelebA-HighCheekbones, as illustrated in E, where high cheekbones is the target label and gender serves as the group attribute.
 Conclusion, limitation, broader impact, and future work
@@
 We will also model the bias between different sensitive groups, e.g., people of some race are less reluctant to reveal their gender.
+Acknowledgement
+We thank the reviewers and the meta-reviewer for assessing our paper and for their constructive feedback. This work is supported by NSF grant RI:1939743.
 Supplementary Materials
@@
 The same approach can be applied to sample $a_{i}$.
-A.4 Soften the step function to enable differentiation
+A.4 ELBO
+We can extend the ELBO as
+$\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ $\geq\!\!\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{z}},a,y)\sim q_{\phi}(\cdot|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})}[\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}|a,y,\mathbf{\bm{z}})\!+\log(p(y)p(a)p(\mathbf{\bm{z}}))-\log q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})]$
+$=\mathop{\mathbb{E}}\limits_{\mathbf{\bm{z}}\sim q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})}\log p_{\theta}(x|a,y,\mathbf{\bm{z}})+\mathop{\mathbb{E}}\limits_{y\sim q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})}\log P_{\theta}({\tilde{y}}|y)+\mathop{\mathbb{E}}\limits_{a\sim q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})}\log P_{\theta}({\tilde{a}}|a)$
+$-\text{KL}[q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})||p(\mathbf{\bm{z}})]-\text{KL}[q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})||p(y)]-\text{KL}[q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})||p(a)]$
+$=:\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}).$
+A.5 Soften the step function to enable differentiation
 The step function [[>0]][\![$\cdot$>0]\!][ [ ⋅ > 0 ] ] offers no useful derivative.
@@
 Given the values of $a_{i}$—either from the training data or from the $N$ samples—it is straightforward to optimize this objective.
-A.5 Labeling test data by minimizing fairness and classification risks
+A.6 Labeling test data by minimizing fairness and classification risks
 With the learned $\phi$, the test data can be labeled by simply seeking the $\mathbf{\bm{y}}$ that maximizes the likelihood while also minimizing the expected fairness risk:
@@
 Afterwards, we do a coordinate descent to finetune $\mathbf{\bm{y}}$ in the face of full expectation on $A_{i}$ for $a_{i}=\emptyset$ using Monte Carlo.
-A.6 Extension to class-probability estimation
+A.7 Extension to class-probability estimation
 The above discussion has been intentionally kept general by considering a randomized classifier for $Y|X$.
@@
 Figure A.2 shows the Pareto frontier of error versus DEO/DEOPP for Adult-Race.
+Figure E shows the results for Celeba-HighCheekbones.
+Here, we use high cheekbones as the label and gender as the group attribute.
 We additionally present the Pareto frontier attained by all the baseline methods.
@@
 ∎
+(a) DEO-Sparse
+(b) DEO-Medium
+(c) DEO-Dense
+(d) DEOPP-Sparse
+(e) DEOPP-Medium
+(f) DEOPP-Dense
+Figure 8: Pareto frontier of error versus DEO/DEOPP for CelebA-HighCheekbones